OpenStudy (anonymous):

I could use some help with a series of questions, anybody up for the challenege? I could use some help with a series of questions, anybody up for the challenege? @Mathematics

OpenStudy (sasogeek):

questions under what topic of math?

OpenStudy (anonymous):

yep

OpenStudy (anonymous):

algebra

OpenStudy (sasogeek):

fair does

OpenStudy (anonymous):

what

OpenStudy (sasogeek):

dont u have the questions?

OpenStudy (anonymous):

forget it, ill ask caleb... ignore this. kinda hope id get help from someone else. no offense.

jimthompson5910 (jim_thompson5910):

what are your questions?

OpenStudy (anonymous):

hey okay so the questions:...pretty easy stuff, i just dunno how i would show my work 50^2+60^2=N x \[\sqrt{625}\] 625

jimthompson5910 (jim_thompson5910):

So you're evaluating 50^2+60^2=N for N?

jimthompson5910 (jim_thompson5910):

Or is it 50^2+60^2=N^2?

OpenStudy (anonymous):

its 50^+60^2=n multiplied by squareroot of 625

jimthompson5910 (jim_thompson5910):

ah ok thx

OpenStudy (anonymous):

mhm.. idk how i would show the steps :/ its one of those things i can just do.

jimthompson5910 (jim_thompson5910):

\[\Large 50^2+60^2 = n\sqrt{625}\] \[\Large 2500+3600 = n(25)\] \[\Large 6100 = 25n\] \[\Large \frac{6100}{25} = n\] \[\Large 244 = n\] \[\Large n = 244\]

OpenStudy (anonymous):

thank you sooo sooo much(:

jimthompson5910 (jim_thompson5910):

np

OpenStudy (anonymous):

i posted another question do you mind helping with it too??